Please help me solve: each of two conguent sides of an isosceles triangle is 8 inches less than twice its base/ the perimeter of the triangle is 74 inches, what is the length of the base?

Answer by rfer(16322)

2x-8+2x-8+x=74
5x-16=74
5x=90
x=18
2x-8=28
28+28+18=74